OpenStudy (cenaida):

Solve the following system of equations using elimination: 3x + 5y = 12 2x + 8y = 22 A. (-1, 3) B. (1, -3) C. (1, 3) D. (3, -1)

OpenStudy (cenaida):

@Luis_Rivera @John_ES @jim_thompson5910

OpenStudy (john_es):

Again, multiply the first row by 2 and the second by -3, \[\begin{cases}6x+10y=24\\ -6x-24y=-66\end{cases}\] Sum the equations \[-14y=-42\Rightarrow y=3\]Now use the substitution in the first equation, for example, and you will obtain x=-1. So you have (-1,3)

OpenStudy (cenaida):

Thanks John!! :D

OpenStudy (john_es):

;)

OpenStudy (john_es):

@Luis_Rivera, you use a wrong system, the y is not in your solution.